OpenStudy (anonymous):

Which of the following is a solution of x² + 14x + 112 = 0? If necessary, round to the nearest hundredth.

OpenStudy (anonymous):

Please help!,

OpenStudy (anonymous):

Try the quadratic formula, \[\frac{-b \pm \sqrt{b^2 - 4ac}}{2a} \] and you will get imaginary numbers as an answer.

OpenStudy (anonymous):

Here is how it should go on your paper. \[ \frac {-14 \pm \sqrt{-252}}{2} = \frac {-14 \pm 6i\sqrt{7}}{2}\] (because 36 x 7 = 252). Now you just have to reduce and you should get \[ 7 \pm 3i \sqrt{7}\]

OpenStudy (anonymous):

oopps.. that should be \[ -7 \pm 3i \sqrt{7}\]